In a U-tube, diameter of two limbs are 0.5 cm and 1 cm respectively. And the tube is filled with water (T= 72 dyne/ cm) then the liquid level difference between the 2 limbs will be

Grade:11

1 Answers

Khimraj
5 years ago
 density of water =   = 1000 kg/m3
radius of limbs , r1 = 0 · 1 cm = 1 × 10-3 m and  r2 = 0 · 5 cm = 5 × 10-3
surface tension , T = 0 · 07 N/m
Now difference in level of water in two tubes ,
h1 - h2 = \frac{2T}{g} (\frac{1}{r_{1}} - \frac{1}{r_{2}})
= 0.01 m 
Hope it clears.........................